OpenStudy (anonymous):

i keep getting 3 and 16 but not sure its correct 5. Solve 3x - 4 = 11 for x. (Points : 1) 3 4 5 6 6. Solve x/4 - 1 = 2 for x. (Points : 1) 4 8 12 16

OpenStudy (anonymous):

3x-4=11 +4 +4 3x=15 /3 /3 x=5 x/4-1=2 +1 +1 x/4=3 *4 *4 x=12

OpenStudy (anonymous):

If you have anymore questions let me know!